You are a German investor and have invested in securities in the Japanese market. Last year, you bought a share of Honda Motor Company on the Tokyo Stock Exchange for Y8500. Recently, you received a dividend of Y250, after which you sold the stock for Y9500. During the year, the exchange rate changed from €0.0091/Y to €0.0083/Y. Compute the return on the Honda stock that you made (as a German investor), and the return that you would have made if you had invested in the Yen, but not in the stock.
Answer #1

Return on stock = (1 + return on stock in foreign currency)*(1 + return on currency) - 1

return on stock in foreign currency = [((ending value or sale value of stock+dividend)/beginning value) - 1]*100

= [((9500+250)/8500) - 1]*100

= [(9750/8500) - 1]*100 = 14.70%

Return on currency = [(current rate/initial rate) - 1]*100

= [(0.0083/0.0091) - 1]*100 = -8.79%

Return on stock = (1+0.1470)*(1-0.0879) - 1 = 4.62%

Return if invested in Yen = [(current exchange rate/initial exchange rate) - 1]*100

= [(0.0083/0.0091) - 1]*100 = -8.79%